1.6 Suppose that a researcher wanted to estimate the parameters in the simple pop- ulation equation Y = βο + β1 X + u. The actual observations on X and Y are INTRODUCTION TO LEAST SQUARES 17 not available, but the following data were calculated from a sample of 10 ob- servations ΣΧΥ= 2.246 Σ xy=-654 у? 41,208 уг-7,568 (a) What is b,? (b) What is bo?
Answer #1

a)

7l7l=b1

=(10*(-654)-(0*0))/(10*310-0^2) = -2.11

b) b0 =

1/n(sum y-b1sum x)

= (1/10)*(0-(-2.11*0))

= 0.1*(0-0) = 0